Calculate the magnitude of Electrostatic force between two human beings
each having mass 100Kg and are 1m apart with 1% excess of electron.
(Molecular mass of H2O= 18gms)
Assume Human Body is made up of 100% water.
1
Expert's answer
2020-02-25T10:24:14-0500
"N=\frac{m}{\mu }N_A=\frac{100}{0.018 }6.02\cdot 10^{23}=3.34\cdot 10^{27}"

"Q=0.01Ne=0.01(3.34\cdot 10^{27})(1.6\cdot 10^{-19})=5.34\cdot 10^{6}C"

"F=k\frac{Q^2}{r^2}"

"F=9\cdot 10^{9}\frac{(5.34\cdot 10^{6})^2}{1^2}=2.6\cdot 10^{23}N"